OpenStudy (anonymous):

SAT: Joan originally had twice as many books as Emily. After she gave Emily 5 books, Joan still had 10 more than Emily. How many books did Joan originally have?

OpenStudy (whpalmer4):

Let J be the number of books Joan originally had. Let E be the number that Emily originally had. J = 2*E After giving 5 books to Emily, Joan still has 10 more books than Emily: J - 5 = E + 5 + 10 You should be able to solve this by substituting "2*E" wherever you see J in that second equation and solving for E. Then use the first equation to find the value of J.